# Perturbation-based Analysis of Compositional Data

### v1

@@ Abstract
Existing statistical methods for compositional data analysis are inadequate for many modern applications for two reasons. First, modern compositional datasets, for example in microbiome research, display traits such as high-dimensionality and sparsity that are poorly modelled with traditional approaches. Second, assessing – in an unbiased way – how summary statistics of a composition (e.g., racial diversity) affect a response variable is not straightforward. In this work, we propose a framework based on hypothetical data perturbations that addresses both issues. Unlike existing methods for compositional data, we do not transform the data and instead use perturbations to define interpretable statistical functionals on the compositions themselves, which we call average perturbation effects. These average perturbation effects, which can be employed in many applications, naturally account for confounding that biases frequently used marginal dependence analyses. We show how average perturbation effects can be estimated efficiently by deriving a perturbation-dependent reparametrization and applying semiparametric estimation techniques. We analyze the proposed estimators empirically on simulated data and demonstrate advantages over existing techniques on US census and microbiome data. For all proposed estimators, we provide confidence intervals with uniform asymptotic coverage guarantees.
Keywords: compositional data, causality, perturbation, feature importance, diversity effects, nonparametrics
@@ 1 Introduction
Compositional data, that is, measurements consisting of parts of a whole, appear in many scientific disciplines, for example, rock compositions in geology, green house gases in the atmosphere, species compositions in ecology and racial or gender distributions in social sciences. The defining property of such data is that they only capture information relative to the unknown whole and hence do not contain any information on the absolute sizes. Depending on the application, the absolute information either might not be of interest (e.g., when analyzing racial or gender distributions) or might not be easily accessible with existing measuring techniques (e.g., in ecology). Mathematically, a $d$-coordinate composition can – without loss of information – be normalized such that it is represented as a point on the $(d-1)$-dimensional simplex
@@
It is well-known that many conventional statistical procedures become invalid when applied to compositional data. For example, already Pearson (1897) drew attention to the fact that correlations between individual compositional components can be spurious due to the sum-to-one constraint. Much work has gone into constructing new procedures for adjusting the correlation measure for compositional components, e.g., by transforming the data with log-ratio transforms (Aitchison, 1982). An instructive viewpoint is to see these approaches not as adjusting the correlation measure but rather as creating new target parameters (e.g., the correlation between log-ratios) that have meaningful interpretations in compositional data.
We adopt this point of view and develop a framework which uses hypothetical perturbations of the data to construct meaningful target parameters. Specifically, we consider a setting where we observe an outcome of interest $Y\in\mathbb{R}$ and an associated compositional predictor $Z\in\Delta^{d-1}$. Our goal is to construct functionals of the conditional distribution $Y$ given $Z$ that are useful when analyzing compositional data. To illustrate this, we consider two examples in which existing approaches are insufficient.
@@
Consider the problem of determining whether the presence or absence of certain microbes in the human gut is affecting a disease. Microbial abundances are commonly measured using genetic sequencing (e.g., targeted amplicon or shotgun metagenomic sequencing) which in most cases only measures the composition and not the absolute abundances. Assume we have measurements from a cohort of patients, where for each patient the relative abundance of all microbes in their gut $Z\in\Delta^{d-1}$ and a disease indicator $Y\in\{0,1\}$ has been recorded. While estimating causal effects in such a setting is often too ambitious, one could hope to screen for relevant microbes that could then be considered for follow-up randomized studies. One possible approach (e.g., Hu et al., 2021) would be to consider a fixed microbe $j$ and then compare the distributions
$Y\,|\,Z^{j}=0\quad\text{and}\quad Y\,|\,Z^{j}\neq 0$
for example, with an independence test or by regressing $Y$ on $\mathbbm{1}_{\{Z^{j}=0\}}$. It is well-known that this type of marginal dependence analysis can be misleading if dependencies exist between the individual components. As an example, assume the disease indicator is generated by
$Y\sim\operatorname{Ber}(\tfrac{3}{4}+\tfrac{1}{8}\mathds{1}_{\{Z^{1}=0\}}-\tfrac{1}{2}\mathds{1}_{\{Z^{2}=0\}}),$
that is, the absence of $Z^{1}$ increases while the absence of $Z^{2}$ decreases the probability of having the disease. In a numerical simulation of this generative model with dependence between $Z^{1}$ and $Z^{2}$, the effect of $Z^{1}$ is instead estimated to decrease the probability of disease as shown in Table 1. If $Z$ did not live on the simplex, we could adjust for the dependence in the remaining coordinates by considering conditional dependencies of the form $Y\mbox{${}\perp\mkern-11.0mu\perp{}$}Z^{j}\,|\,Z^{-j}$. However, due to the simplex constraint $Z^{-j}$ completely determines $Z^{j}$, so the conditional independence always holds. In the Table 1, we also provide the results from applying the compositional knockout effect to $Z^{1}$ $(\mathrm{CKE}^{1})$ which we introduce in Section 2.2.
@@
OLS of $Y$ on $\mathbbm{1}_{\{Z^{1}=0\}}$ $-0.12$ $(-0.17,-0.06)$
$\mathrm{CKE}^{1}$ using rf $0.15$ $(0.10,0.21)$
@@
OLS of Y on $-G(Z)$ $-1.57$ $(-2.18,-0.97)$
$\mathrm{CDI}_{\mathrm{Gini}}$ using rf $0.63$ $(0.14,1.12)$
Table 1: Numerical illustration of Examples 1 and 2. The first row in each table corresponds to the marginal effect estimate based on a correctly specified linear model. The second row correspond to estimates of two of the target parameters introduced in Section 2.2 with random forest (rf) as nuisance function estimators. (Left) The estimated marginal effect is negative, even though the true effect is positive. The compositional knock-out effect ($\mathrm{CKE}^{1}$), which adjust for the simplex structure, results in the correct sign. (Right) The estimated marginal effect of an increase in diversity (measured by the negative Gini coefficient, i.e., $-G(Z)$) is negative even though the true effect is equal to $1$ by construction. The compositional diversity effect $\mathrm{CDI}_{\mathrm{Gini}}$ correctly recovers the effect. Details on both experiments are provided in Section S4.1 of the supplementary material.
@@
Consider the problem of determining the effect of an increase in diversity of $Z$ on the response $Y$. Such a question is of immediate interest in many applied fields, e.g. when investigating whether more racially diverse university programs improve educational outcomes (Antonio et al., 2004) or whether increased income equality boosts economic output (Shin, 2012). The prevalent approach is to summarize diversity via a one-dimensional summary statistic $G(Z)$ (e.g., the Gini coefficient, generalized entropy or Gini-Simpson index) and analyze the correlation between $G(Z)$ and $Y$. As in Example 1 this type of marginal dependence analysis is potentially misleading in practice. In Table 1 we illustrate this on a specific numeric example, where the true causal effect is positive, but the marginal method leads to a negative effect of diversity (details are provided in Section S4.1 of the supplementary material). To avoid this, one would again like to adjust for the remaining dependence within $Z$ to avoid the confounding biases. However, since $G(Z)$ depends on all of $Z$ existing adjustment approaches are not applicable. The compositional diversity effect (with Gini-speed) which we introduce in Section 2.2 appropriately adjust for the confounding as shown in Table 1. A more in-depth example in which the marginal approach is misleading is given in Section 4.2 where we estimate the effect of racial diversity on income on grouped real data. Unlike in Example 1, the problem of not being able to adjust with existing techniques is not solely due to the simplex-constraint and similarly appears for summary statistics of unconstrained Euclidean data. Our perturbation-based framework can also be extended to those settings, but as the use of this type of analysis is particularly prevalent in compositional data applications we restrict our attention to this setting.
In this work, we propose a procedure that accounts for potential confounding within $Z$. The main idea is to introduce a new target parameter via a hypothetical perturbation of the data. More concretely, a perturbation is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a path $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed. For example, in Example 1 when considering the first microbe (i.e., $j=1$), we could consider a binary perturbation $\psi$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one. A target parameter that captures the desired adjusted effect is then given by
$\lambda_{\psi}:=\frac{\mathbb{E}\left[\mathbb{E}[Y\,|\,Z=\psi(Z,1)]\right]-\mathbb{E}\left[\mathbb{E}[Y\,|\,Z=\psi(Z,0)]\right]}{\mathbb{P}(Z\neq\psi(Z,1))}.$
@@
In Example 2, an increase in the diversity of $Z$ can be expressed as a perturbation $\psi$ that pushes every point in $\Delta^{d-1}$ to a more diverse point. Formally, using the center of the simplex (i.e., $z_{\mathrm{cen}}=(1/d,\ldots,1/d)\in\Delta^{d-1}$) as the most diverse point, we define for all $z\in\Delta^{d-1}$ and all $\gamma\in[0,\|z_{\mathrm{cen}}-z\|_{1}]$ the perturbation
@@
We can then interpret $\psi(z,\epsilon)$ for a small $\epsilon>0$ as slightly increasing the diversity of the point $z$. With this notion of an increase in diversity, a target parameter that captures the desired adjusted effect is given by
$\tau_{\psi}:=\mathbb{E}\left[\partial_{\gamma}\mathbb{E}\left[Y\,|\,Z=\psi(Z,\gamma)\right]\big{|}_{\gamma=0}\right],$
@@
While the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ may seem simple at first sight, they can be challenging to estimate directly as they depend on estimating the conditional expectation $z\mapsto\mathbb{E}[Y\,|\,Z=z]$ which due the simplex structure is generally a complex functional. We therefore propose an approach that reparametrizes the compositional vector $Z\in\Delta^{d-1}$ via a bijection $\phi$ to a vector $(L,W)\in\mathbb{R}\times\mathcal{W}$ in such a way that the perturbation only changes $L$ in the new parametrization. To make this more concrete, assume that the structural causal model (SCM) (e.g., Pearl, 2009) given in Figure 1 (left) generates the data (the causal terminology is used only to provide a more intuitive introduction). Given a perturbation $\psi$ that we are interested in, we now propose to construct a bijection $\phi:\Delta^{d-1}\rightarrow\mathbb{R}\times\mathcal{W}$ that maps $Z$ to $(L,W)$ such that if we intervene on $Z$ by setting it to $\psi(Z,\gamma)$ this corresponds to an intervention in an SCM over the variables $(Y,L,W)$ on $L$ alone (see Figure 1 (right)). Hence, if we are able to construct such a reparametrization, then we have successfully reduced our problem to a setting in which standard adjustment works, that is, using any existing method to estimate the effect of $L$ on $Y$ that adjusts for $W$ results in an unbiased estimate of the causal effect. Given a perturbation $\psi$, we provide a general recipe to construct the reparametrization $\phi$. We furthermore propose several explicit perturbations (including their corresponding reparametrizations) and suggest efficient estimators of the perturbation-based target parameters using semiparametric theory based on the $(L,W)$-reparametrization.
@@
$L$ $\leftarrow m(W,\xi)$ $Y$ $\leftarrow h(L,W,\varepsilon)$

### v2

# Perturbation-based Effect Measures for Compositional Data

@@ Abstract
Existing effect measures for compositional features are inadequate for many modern applications for two reasons. First, modern datasets with compositional covariates, for example in microbiome research, display traits such as high-dimensionality and sparsity that can be poorly modelled with traditional parametric approaches. Second, assessing – in an unbiased way – how summary statistics of a composition (e.g., racial diversity) affect a response variable is not straightforward. In this work, we propose a framework based on hypothetical data perturbations that addresses both issues. Unlike many existing effect measures for compositional features, we do not define our effects based on a parametric model or a transformation of the data. Instead, we use perturbations to define interpretable statistical functionals on the compositions themselves, which we call average perturbation effects. These effects naturally account for confounding that biases frequently used marginal dependence analyses. We show how average perturbation effects can be estimated efficiently by deriving a perturbation-dependent reparametrization and applying semiparametric estimation techniques. We analyze the proposed estimators empirically on simulated and semi-synthetic data and demonstrate advantages over existing techniques on data from New York schools and microbiome data. For all proposed estimators, we provide confidence intervals with uniform asymptotic coverage guarantees.
Keywords: compositional data, effect estimation, causality, perturbation, feature importance, diversity effects, nonparametrics, mixture experiments
@@ 1 Introduction
Compositional data, that is, measurements consisting of parts of a whole, appear in many scientific disciplines, for example, rock compositions in geology, green house gases in the atmosphere, species compositions in ecology and racial or gender distributions in social sciences. The defining property of such data is that they only capture information relative to the whole. Depending on the application, the absolute information either might not be of interest (e.g., when analyzing racial or gender distributions) or might not be easily accessible with existing measuring techniques (e.g., in ecology). Mathematically, a $d$-coordinate composition can – without loss of information – be normalized such that it is represented as a point on the $(d-1)$-dimensional simplex
@@
Modelling the relationship between a real-valued response $Y\in\mathbb{R}$ and a composition $Z\in\Delta^{d-1}$ has a long history and has, for example, been considered in the literature on ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002) and ’log-ratio based compositional data analysis’ (Aitchison, 1982), which we discuss in more detail in Section 1.1. A key aspect of this research is that it attempts to achieve two separate goals simultaneously: (i) ensuring that the model parameters are interpretable and (ii) providing parametric models that fit to realistic data-generating processes with compositional covariates. More recently, in the context of high-dimensional and sparse data, alternative approaches (Knight et al., 2018; Cammarota et al., 2020; Huang et al., 2023) to solving the second goal have been proposed that apply flexible machine learning methods to model the relationship between $Y$ and $Z$ nonparametrically. However, such approaches are at odds with the first goal of providing interpretable effect measures and cannot be combined with the above-mentioned approaches which heavily rely on parametric models to define such effects.
Our main contribution is to provide a unified framework for defining interpretable effect measures for compositional features which we call average perturbation effects that can be combined with flexible nonparametric estimation procedures. Special cases of these average perturbation effects have interpretations that are similar to some model-based individual feature effect measures, e.g., Cox-directions (Cox, 1971) and log-contrast coefficients (Aitchison and Bacon-Shone, 1984). Average perturbation effects, however, are defined without reference to a specific regression model and therefore do not rely on a correctly specified parametric model. Importantly, this allows us to construct more flexible target quantities, two of which we believe are particularly relevant for applications. First, effects based on binary perturbations are useful when summarizing the effect of a discrete change, e.g. the effect of setting $Z^{j}$ to 00 while rescaling $Z^{-j}$ to the simplex. In Section 4.3 we show empirically that the compositional features with significant effects as measured by these binary perturbations coincide with those deemed important by the standard log-contrast regression procedures used in the microbiome literature. Second, effects based on directional perturbations summarize continuous changes in specific directions. This can involve changing individual components of $Z$, but also allows for more complex changes that are not tied to a specific coordinate of $Z$. For example, we can define the effect of ‘diversifying’ $Z$, i.e. moving closer to the center of $\Delta^{d-1}$. Such diversifying effects have been considered previously (Antonio et al., 2004) but are usually defined as the effect of marginally changing a specific measure of diversity. However, as the effect of diversifying can be confounded by other features of $Z$ such a marginal analysis can be misleading and due to the simplex constraint, there is no immediate way to correct for this confounding when estimating these effects (see Example 2 and Section 4.2). In contrast, our perturbation-based framework can control for the remaining variation in $Z$ leading to unconfounded effect estimates with valid asymptotic inference.
Our proposal is intimately related to the growing literature on targeted learning and semiparametric inference (Chernozhukov et al., 2018; Berk et al., 2021; Vansteelandt and Dukes, 2022). An important aspect of our work is, however, to clarify the need for tailored target estimands in the presence of compositional structure in the covariates. Applying a nonparametric effect measure that is not aware of the sum-to-one constraint of the compositional vector can be just as misleading as naively estimating a non-compositional effect based on a parametric model (see Example 1 and Section 4.3). Beyond defining average perturbation effects, we also provide several algorithm-agnostic procedures to estimate the average perturbation effects. These procedures are essentially well-known semiparametric learning algorithms applied after appropriately transforming the compositional covariates. While these algorithms require us to estimate nuisance functions relating to the distribution of $(Y,Z)$, which are particularly delicate for compositional covariates, they only require these estimates to be accurate and not interpretable. This allows us to apply arbitrary regression methods as long as they are sufficiently predictive.
Average perturbation effects are defined by comparing the change in the expected response of $Y$ under a hypothetical perturbation of the composition $Z$. To motivate this, consider Euclidean covariates $X\in\mathbb{R}^{d}$ instead of the composition $Z\in\Delta^{d-1}$. Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{d})\in\mathbb{R}^{d}$ such that for all $x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}$, $f(x):=\mathbb{E}[Y\,|\,X=x]=\sum_{j=1}^{d}\beta^{j}x^{j}$, that is, the conditional expectation of $Y$ given $X=x$ is a linear function. Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the rate of change in the $j$th coordinate, or more formally,
$\partial_{x_{j}}f(x^{1},\dots,x^{d})=\beta^{j}.$
We can rephrase this interpretation by considering a perturbation $\psi:\mathbb{R}^{d}\times[0,1)\rightarrow\mathbb{R}^{d}$ that additively increases the $j$th coordinate, that is, $\psi(x,\gamma)=(x^{1},\dots,x^{j}+\gamma,\dots,x^{d})$. Then, for any $x\in\mathbb{R}^{d}$, we have
$\partial_{x_{j}}f(x)=\lim_{\gamma\to 0}\frac{f(\psi(x,\gamma))-f(\psi(x,0))}{\gamma}=\partial_{\gamma}f(\psi(x,\gamma))\bigm{|}_{\gamma=0}.$
Therefore, the coefficient $\beta^{j}$ also corresponds to the initial rate of change (initial referring to $\gamma=0$) of $f$ along the perturbation $\psi$. Similar observations are possible regarding other interpretable effect measures for Euclidean covariates and under more general nonlinear models.
Returning now to the composition $Z$ and redefining for $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$, $f(z):=\mathbb{E}[Y\,|\,Z=z]$, we see that neither of the above expressions are well-defined. The quantity $f(\psi(z,\gamma))$ is meaningless for any non-zero $\gamma\in\mathbb{R}$, since $f$ is only defined on $\Delta^{d-1}$. We can however use similar ideas by defining tailored perturbations on $\Delta^{d-1}$ to obtain interpretable effect measures. The following two examples provide toy applications that illustrate the pitfall and short-comings of existing approaches.
@@
Consider the problem of determining whether the presence or absence of certain microbes in the human gut is affecting a disease. Microbial abundances are commonly measured using genetic sequencing (e.g., targeted amplicon or shotgun metagenomic sequencing) which in most cases only measures the composition and not the absolute abundances. Assume we have measurements from a cohort of patients, where for each patient the relative abundance of all microbes in their gut $Z\in\Delta^{d-1}$ and a disease indicator $Y\in\{0,1\}$ ($1$ encoding disease) has been recorded. While estimating causal effects in such a setting is often too ambitious, one could hope to screen for relevant microbes that could then be considered for follow-up randomized studies. One possible approach (e.g., Hu et al., 2021) would be to consider a fixed microbe $j$, define the treatment variable $T:=\mathds{1}_{\{Z^{j}=0\}}$ corresponding to the presence or absence of microbe $j$ and then estimate a marginal treatment effect given by
$\mathrm{MTE}:=\mathbb{E}[Y\,|\,T=1]-\mathbb{E}[Y\,|\,T=0]=\mathbb{P}(Y=1\,|\,T=1)-\mathbb{P}(Y=1\,|\,T=0),$
where we implicitly assume $\mathbb{P}(T=0)>0$ and $\mathbb{P}(T=1)>0$. It is well-known that this type of marginal dependence analysis can be misleading if dependencies exist between the individual components. [^1] This is true when $Z$ is Euclidean, but the problem is even more severe when $Z$ is compositional as the sum-to-one constraint enforces additional dependency between the components of $Z$. To illustrate this, assume the disease indicator given the composition has conditional distribution for all $z\in\Delta^{d-1}$ given by
$Y\,|\,Z=z\sim\operatorname{Ber}(\tfrac{3}{4}+\tfrac{1}{8}\mathds{1}_{\{z^{1}=0\}}-\tfrac{1}{2}\mathds{1}_{\{z^{2}=0\}}),$
that is, the absence of $Z^{1}$ increases while the absence of $Z^{2}$ decreases the probability of having the disease. In a numerical simulation of this generative model with dependence between $Z^{1}$ and $Z^{2}$, the effect of $Z^{1}$ is instead estimated to decrease the probability of disease as shown in Table 1. If $Z$ was Euclidean, we could adjust for the dependence in the remaining coordinates by estimating a conditional treatment effect such as
$h:z^{-j}\mapsto\mathbb{E}[Y\,|\,T=1,Z^{-j}=z^{-j}]-\mathbb{E}[Y\,|\,T=0,Z^{-j}=z^{-j}]$
and marginalizing it to produce a single effect measure $\mathbb{E}[h(Z^{-j})]$. [^2] In the literature on causal inference (and under appropriate causal assumptions), $h$ corresponds to the conditional average treatment effect (CATE). Estimating this quantity is a well-studied problem, e.g. using the AIPW estimator (Robins et al., 1994) or using a partially linear model (Chernozhukov et al., 2018). Unfortunately, neither of these estimators are well-defined when $Z$ is compositional as it is not possible to change $T$ and $Z^{-j}$ independently of each other and, in fact, for all $z^{-j}$, we have $h(z^{-j})=0$ as $Y\mbox{${}\perp\mkern-11.0mu\perp{}$}T\,|\,Z^{-j}$ due to the sum-to-one constraint which makes $T$ fully determined by $Z^{-j}$. In the Table 1, we provide the results from applying our proposed marginalized conditional treatment effect the compositional knockout effect to $Z^{1}$ $(\mathrm{CKE}^{1})$ which we introduce in Section 2.2.
@@
$\mathrm{MTE}$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\mathbf{-0.12}}$ $(-0.17,-0.06)$
$\mathrm{CKE}^{1}$ using rf ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.15}}$ $(0.10,0.21)$
@@
OLS of $Y$ on $-G(Z)$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\mathbf{-1.57}}$ $(-2.18,-0.97)$
$\mathrm{CDI}_{\mathrm{Gini}}$ using rf ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.63}}$ $(0.14,1.12)$
Table 1: Numerical illustration of Examples 1 and 2. The first row in each table corresponds to the marginal treatment effect (MTE or OLS of $Y$ on $-G(Z)$) estimate based on a correctly specified linear model. The second row correspond to estimates of two of the target parameters introduced in Section 2.2 with random forest (rf) as nuisance function estimators. Bold red numbers indicate a significant negative effect at a $5\%$ level, while bold green indicates a significant positive effect. (Left) The estimated MTE is negative, even though the true effect is positive. The compositional knock-out effect ($\mathrm{CKE}^{1}$), which adjust for the simplex structure, results in the correct sign. (Right) The estimated marginal effect of an increase in diversity (measured by the negative Gini coefficient, i.e., $-G(Z)$) is negative even though the true effect is equal to $1$ by construction. The compositional diversity effect $\mathrm{CDI}_{\mathrm{Gini}}$ correctly recovers the effect. Details on both experiments are provided in Section S4.1 of the supplementary material.
@@
Consider the problem of determining the effect of an increase in diversity of $Z$ on a response $Y$. Such a question is of immediate interest in many applied fields, e.g. when investigating whether more racially diverse university programs improve educational outcomes (Antonio et al., 2004) or whether increased income equality boosts economic output (Shin, 2012). The prevalent approach is to summarize diversity via a one-dimensional summary statistic $G(Z)$ (e.g., the Gini coefficient, generalized entropy or Gini-Simpson index) and analyze the correlation between $G(Z)$ and $Y$. As in Example 1 this type of marginal dependence analysis is potentially misleading in practice. In Table 1 we illustrate this on a specific numeric example, where the true causal effect is positive, but the marginal method leads to a negative effect of diversity (details are provided in Section S4.1 of the supplementary material). To avoid this type of confounding bias, one would again like to adjust for the remaining dependence within $Z$. However, since $G(Z)$ depends on all of $Z$ existing adjustment approaches are not applicable. The compositional diversity effect (with Gini-speed) which we introduce in Section 2.2 appropriately adjusts for the confounding as shown in Table 1. A more in-depth example in which the marginal approach is misleading is given in Section 4.2.1 where we estimate the effect of racial diversity on income on a semi-synthetic dataset based on US census data. A real application is given in Section 4.2.2 where the effect of racial diversity on school grades is analyzed using data from schools in New York. Unlike in Example 1, the problem of not being able to adjust with existing techniques is not solely due to the simplex-constraint and similarly appears for summary statistics of unconstrained Euclidean data. Our perturbation-based framework can also be extended to those settings, but as the use of this type of analysis is particularly prevalent in compositional data applications we restrict our attention to this setting.
We postpone the formal definitions of perturbations and average perturbation effects to Section 2 but give a brief description here. A perturbation is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a trajectory $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed. For example, in Example 1 when considering the first microbe (i.e., $j=1$), we could consider a binary perturbation $\psi:\Delta^{d-1}\times\{0,1\}\rightarrow\Delta^{d-1}$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one. By defining $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ as the regression function, we can then consider contrasts of the form $f(\psi(z,1))-f(\psi(z,0))$, which capture the effect of setting $z^{1}=0$ while the other coordinates are proportionally rescaled to account for the change in $z^{1}$. The fact that $Z$ is compositional necessitates that we fully specify what happens to all of $Z$ rather than only modifying $Z^{j}$ as one would do in the usual Euclidean approach outlined in Example 1. By marginalizing over $Z$ the perturbation-based approach leads to the average effect
$\mathbb{E}[f(\psi(Z,1))-f(\psi(Z,0))].$
In order to make the effect measure agnostic to the proportion of $Z$s that already have $Z^{j}=0$ in the population, we instead propose to consider the effect measure
$\lambda_{\psi}:=\frac{\mathbb{E}[f(\psi(Z,1))-f(\psi(Z,0))]}{\mathbb{P}(Z\neq\psi(Z,1))}.$
@@
$\lambda_{\psi}$ measures the expected change in $Y$ when setting $Z^{j}=0$ for an individual that does not already have $Z^{j}=0$. In Example 2, an increase in the diversity of $Z$ can be expressed as a perturbation $\psi$ that pushes every point in $\Delta^{d-1}$ to a more diverse point. Formally, using the center of the simplex (i.e., $z_{\mathrm{cen}}=(1/d,\ldots,1/d)\in\Delta^{d-1}$) as the most diverse point, we define for all $z\in\Delta^{d-1}$ and all $\gamma\in[0,\|z_{\mathrm{cen}}-z\|_{1}]$ the perturbation
@@
We can then interpret $\psi(z,\epsilon)$ for a small $\epsilon>0$ as slightly increasing the diversity of the point $z$. With this notion of an increase in diversity, we propose to measure the effect of increasing diversity by
$\tau_{\psi}:=\mathbb{E}\left[\partial_{\gamma}f(\psi(Z,\gamma))\big{|}_{\gamma=0}\right],$
@@
While the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ may seem simple at first sight, they can be challenging to estimate directly as they depend on estimating the regression function $z\mapsto\mathbb{E}[Y\,|\,Z=z]$ which, in part due to the simplex structure, is generally a complex functional. We therefore propose an approach that reparametrizes the compositional vector $Z\in\Delta^{d-1}$ via a bijection $\phi$ to a vector $(L,W)\in\mathbb{R}\times\mathcal{W}$ in such a way that the perturbation only changes $L$ in the new parametrization. To make this more concrete, assume that the structural causal model (SCM) (e.g., Pearl, 2009) given in Figure 1 (left) generates the data (the causal terminology is used only to provide a more intuitive introduction). Given a perturbation $\psi$ that we are interested in, we now propose to construct a bijection $\phi:\Delta^{d-1}\rightarrow\mathbb{R}\times\mathcal{W}$ that maps $Z$ to $(L,W)$ such that if we intervene on $Z$ by setting it to $\psi(Z,\gamma)$ this corresponds to an intervention in an SCM over the variables $(Y,L,W)$ on $L$ alone (see Figure 1 (right)). Hence, if we are able to construct such a reparametrization, then we have successfully reduced our problem to a setting in which standard adjustment works, that is, using any existing method to estimate the effect of $L$ on $Y$ that adjusts for $W$ results in an unbiased estimate of the causal effect. Given a perturbation $\psi$, we provide a general recipe to construct the reparametrization $\phi$. We furthermore propose several explicit perturbations (including their corresponding reparametrizations) and suggest efficient estimators of the perturbation-based target parameters using semiparametric theory based on the $(L,W)$-reparametrization.
@@
$L$ $\leftarrow m(W,\xi)$
$Y$ $\leftarrow h(L,W,\varepsilon)$